OpenStudy (jennyrlz):

Write the sum using summation notation, assuming the suggested pattern continues. -10 - 2 + 6 + 14 + ... + 110

OpenStudy (michele_laino):

your sequence is an arithmetic sequence with constant d= 8

OpenStudy (michele_laino):

first term is: \[{a_1} = - 10\]

OpenStudy (jennyrlz):

ok i understand that

OpenStudy (michele_laino):

n-th term is: \[{a_n} = 110\]

OpenStudy (michele_laino):

what is n?

OpenStudy (michele_laino):

or, in other words how many terms have you got in your sequence?

OpenStudy (jennyrlz):

5 terms?

OpenStudy (jennyrlz):

so n is equal to 5

OpenStudy (michele_laino):

I don't think, you have to apply this formula: \[{a_n} = {a_1} + (n - 1) \cdot d\] where d= 8, a_1 = -10, and a_n = 110, so please solve that equation for n

OpenStudy (jennyrlz):

ok one sec

OpenStudy (michele_laino):

hint: \[n - 1 = \frac{{{a_n} - {a_1}}}{d} = \frac{{110 - \left( { - 10} \right)}}{8} = ...?\]

OpenStudy (jennyrlz):

15 :D

OpenStudy (michele_laino):

please wait you got: n-1 = 15 so what is n?

OpenStudy (jennyrlz):

14

OpenStudy (solomonzelman):

just find the pattern, and put it in the \(\large\color{blueviolet}{ \displaystyle \sum_{ }^{ } ~}\)

OpenStudy (solomonzelman):

yes, find whichTH term 110 is, and do the sigma

OpenStudy (michele_laino):

hint: if n-1 = 15, then n = 15+1=...?

OpenStudy (solomonzelman):

if anyone needs the latex for inline sigma, then ``` \(\large\color{blueviolet}{ \displaystyle \sum_{ n=1 }^{ \infty } ~ f(n)}\) ```

OpenStudy (jennyrlz):

N=16

OpenStudy (michele_laino):

that's right! so in you sequence there are 16 terms

OpenStudy (jennyrlz):

Ok!

OpenStudy (michele_laino):

now, in order to get the sum S of tose terms, you have to apply this formula: \[S = \frac{{{a_1} + {a_n}}}{2}n\]

OpenStudy (michele_laino):

oops.. those terms...

OpenStudy (michele_laino):

hint: \[S = \frac{{{a_1} + {a_n}}}{2}n = \frac{{ - 10 + 110}}{2} \cdot 16 = ...?\]

OpenStudy (jennyrlz):

800...?

OpenStudy (michele_laino):

well done! Congratulations! :)

OpenStudy (jennyrlz):

:D

OpenStudy (jennyrlz):

Is that it...?

OpenStudy (michele_laino):

yes! it is!

OpenStudy (jennyrlz):

But it doesn't match any of the answers given 0.0

OpenStudy (jennyrlz):

The answers have the "E"

OpenStudy (jennyrlz):

I can't make it on this key board :/

OpenStudy (michele_laino):

please wait I think that @SolomonZelman is right! since we have to write in the sum notation

OpenStudy (jennyrlz):

Ok :)

OpenStudy (michele_laino):

so our result, since there are infinite terms is: \[\sum\limits_1^\infty {{a_1} + \left( {n - 1} \right)} \,d\]

OpenStudy (jennyrlz):

S there are no longer 16 terms...?

OpenStudy (michele_laino):

please you have to substitute a_1 and d with our numerical values

OpenStudy (jennyrlz):

A1=-10 d=8

OpenStudy (michele_laino):

no I don't think, since your problem says "assuming that the pattern"

OpenStudy (michele_laino):

ok! sbstitute a_1 = -10 andd = 8, into this formula \[\sum\limits_1^\infty {{a_1} + \left( {n - 1} \right)} \,d\]

OpenStudy (jennyrlz):

Ahh ok

OpenStudy (jennyrlz):

So what do I put for n?

OpenStudy (jennyrlz):

E(-10)+(8n-8)

OpenStudy (michele_laino):

please unchanged n, write n without any substitution for it

OpenStudy (michele_laino):

yes! now simplify like terms, namely what is: -10+8n-8=...?

OpenStudy (jennyrlz):

?

OpenStudy (jennyrlz):

So (n-8)?

OpenStudy (michele_laino):

you should get this: \[\sum\limits_1^\infty {8n - 18} \]

OpenStudy (jennyrlz):

-18

OpenStudy (michele_laino):

since -10 + 8n - 8= 8n-18

OpenStudy (michele_laino):

no, 8n-18

OpenStudy (jennyrlz):

I am on an I pad so sorry if everything is taking forever to load if for you

OpenStudy (michele_laino):

please we have: (-10)+(8n-8)= 8n-18

OpenStudy (michele_laino):

+ 8n

OpenStudy (michele_laino):

-18 + 8n

OpenStudy (michele_laino):

(-10)+(8n-8) = -10 + 8n -8 = -10 - 8 +8n = -18 + 8n

OpenStudy (jennyrlz):

That still doesn't match any of my answers T.T

OpenStudy (michele_laino):

it is a simplification, since I have combined similar terms

OpenStudy (jennyrlz):

E=-80n E=(-10+8n)

OpenStudy (jennyrlz):

Tose are the answer choices

OpenStudy (michele_laino):

and the result is: \[\sum\limits_1^\infty {8n - 18} \]

OpenStudy (jennyrlz):

those done match any of my answer options

OpenStudy (michele_laino):

please is it clear?

OpenStudy (jennyrlz):

E=-80n or E=(-10+8n)

OpenStudy (michele_laino):

if not I will give you more explanation

OpenStudy (jennyrlz):

are you still there? ;-;

OpenStudy (jennyrlz):

wait you lost me ; ;

OpenStudy (jennyrlz):

the answer we got doesnt match any of the answer choices

OpenStudy (michele_laino):

ok! I'm here!

OpenStudy (michele_laino):

please what are your choices?

OpenStudy (jennyrlz):

ok the answer choices are either E=-80n or E= (-10+8n)

OpenStudy (michele_laino):

then it is the second option

OpenStudy (jennyrlz):

ok and there are two of the "second" option on has a 15 over the E and the other has an infinity symbol over it

OpenStudy (michele_laino):

please wait, are you reading them from your textbook?

OpenStudy (jennyrlz):

no

OpenStudy (jennyrlz):

this is from an assignment the teacher gave us

OpenStudy (michele_laino):

then you have to pick the one which has an infinity symbol

OpenStudy (jennyrlz):

ok thanks :DDD

OpenStudy (michele_laino):

thanks! :)